OpenStudy (anonymous):

find dy/dx when y=0 if x cosy - sinx - 2 = 0

OpenStudy (loser66):

1/ isolate y 2/ take differential of y 3/ dat sit

OpenStudy (anonymous):

would you mind showing me how to isolate y? im not sure how to do that

OpenStudy (anonymous):

differentiate w.r.t. x \[x(-\sin y)\frac{ dy }{ dx }+1*\cos y-\cos x-0=0\] \[\frac{ dy }{dx }=?\]

OpenStudy (anonymous):

\[\frac{ -\cos x }{ x \sin y }\]

OpenStudy (anonymous):

\[\frac{ dy }{dx }=\frac{ \cos y-\cos x }{x \sin y }\]

OpenStudy (loser66):

Let me try once xcos y - sinx -2 =0 xcos y = 2 + sinx cos y = \(\dfrac{2+sinx}{x}\) y = arcos \(\dfrac{2+sinx}{x}\) now, take derivative the right hand side